8
_ + 1

(2/x)+ 1
1
Expert's answer
2011-02-24T05:54:45-0500
8/(2/x + 1) +1 = 8x/(2+x) +1 = (8x + 2 +x) /(2 +x) = (9x +2)/(2+x)